P(A U B) = 15/16; P(A ∩ B) = 1/8; P(A) =3/8. Find P(B).

Grade:

1 Answers

vijesh jayan
13 years ago

P(A U B) = P(A) + P(B) - P(A ∩ B)

15/16 = 3/8 + P(B) - 1/8

15/16 = 2/8 + P(B)

P(B) = 15/16 - 2/8 = (15 - 4)/16 = 11/16

if right approve plzzzzzzzzzzzzzzzzzzzzz